A deposit of N100 is made at the beginning of each month in an account that pays 6% interest, compounded monthly. The balance in the account at the end of 5 years is? 


1
Expert's answer
2021-12-17T13:39:55-0500

A deposit of $100 is made at the beginning of each month in an account that pays 6% interest, compounded monthly. The balance A in the account at the end of 5 years is

"A=100(1 +\frac{0.06}{12})^1 +...+100(1 +\frac{0.06}{12})^{60}"

The balance A in account at the end of 5 years is,

"A=100(1 +\frac{0.06}{12})^1 + 100(1 + \frac{0.06}{12})^2+...+100(1 +\frac{0.06}{12})^{60}"

We can see that,

"a_1=100(1+\frac{0.06}{12}) \\

r=(1 +\frac{0.06}{12}) \\

n=60"

Now, we can apply the formula for the sum of a finite geometric sequence,

"A=S_n \\

=a(\frac{1-r^n}{1-r}) \\

=100(1+\frac{0.06}{12})(\frac{1-1.005^{60}}{1-1.005}) \\

= 7011.89"

Therefore the balance A in account at the end of 5 years is 7011.89